# African Data Ethics: A Discursive Framework for Black Decolonial Data Science

### v1

@@ 4. African Data Science Ethical Framework
In this section, we summarize the key components of our framework, shown in Table 1.
Major Principle | Minor Principles |
Challenge Colonial Power | |
Decolonize & Challenge Internal Power Asymmetry | Challenge Internal Power Asymmetry |
Community in Everything | |
Solidarity | |
Inclusion of the Marginalized | |
Center Remote & Rural Communities | |
Center All Communities | Center Women |
Universal Dignity | |
Common Good | |
Uphold Universal Good | Harmony |
Consensus-Building | |
Reciprocity | |
Resolving Data-Driven Harms | |
Communalism in Practice | Fair Collaboration |
”For Africans, By Africans” | |
Treasure Indigenous Knowledge | |
Data Self-Determination | Data Sovereignty & Privacy |
Measured Development | |
Technical Infrastructure | |
Governance Infrastructure | |
Invest in Data Institution & Infrastructures | Support Formal & Informal Collectives |
Holistic Education | |
Prioritize Education & Youth | Youth Empowerment and Protection |
4.1. Decolonize & Challenge Internal Power Asymmetry
@@
Challenge Colonial Power. RDS practices from the West do not seamlessly transfer to the African context because these practices are developed within colonial contexts disconnected from the realities of African practitioners and users (Eke and Ogoh, 2022; Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023; Adelani et al., 2022; Gwagwa, 2019; Eke et al., 2023b; Okolo et al., 2023; Eke et al., 2023a; Goffi, 2023; Carman and Rosman, 2023). African practitioners identify three dimensions in which colonialism and imperialism limit RDS: epistemic injustice, dehumanizing extraction, and dependent partnerships. Firstly, African scholars identify trends in philosophical epistemic injustice permeating global data ethics paradigms (Eke and Ogoh, 2022; Metz, 2021; Olojede, 2023). As many African philosophers agree, Enlightenment ideals (a premier part of the Western philosophical canon) were predicated on colonialism and racism (Gwagwa, 2019). Africans were deemed incapable of rational thinking by Western colonizers, so through the Enlightenment principle of rationality, anti-blackness was justified (Lauer, 2017). Furthermore, colonization was not only excused but encouraged by rationality so colonizers could develop Africans through Western instruction. Under colonial rule, Africans were taught to abandon their Indigenous knowledge to adopt the rational knowledge of the West. The legacy of colonialism is why African data scientists encourage casting aside Western perspectives to develop African RDS perspectives (Mhlambi, 2020). Additionally, an over reliance on performance metrics encourages the same colonial blindspot that excuses and encourages the marginalization of Africans in technology such as facial recognition (Mhlambi, 2020; Buolamwini and Gebru, 2018; Cisse, 2018; Gwagwa et al., 2022).
@@
Secondly, many documents recognize that most African contributions to data science disproportionately benefit Western corporations like OpenAI, Google, Meta, and Microsoft (Ndjungu, 2020; Chan et al., 2021; Abebe et al., 2021; Nwankwo and Sonna, 2019; Kiemde and Kora, 2022). The computing demand of large-data systems such as AI proliferates neocolonialism to new heights in Africa (Eke et al., 2023a). The work of Africans within the data science ecosystem should benefit Africans first (Kohnert, 2022). The fact that it currently does not is connected to the legacy of colonialism and chattel slavery in which Africans were forced to extract their raw materials so colonialists could fuel industrialization and capitalism in their home countries (Mhlambi, 2020; Ndjungu, 2020; Day and Rennie, 2023; Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023; Birhane, 2020).
@@
Finally, the last vestige of colonial power to be challenged in African RDS is dependent partnership. Africa currently lacks the technical infrastructure for large-scale DDS, which pushes data scientists towards unfair agreements with powerful organizations to gain technology (Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023; Hountondji, 2004; Osaghae, 2004). Even worse, companies such as Amazon, Google, Meta, and Uber use savior language such as “liberating the bottom million” to describe their digital services in Africa (Abebe et al., 2021).
@@
Challenge Internal Power Asymmetry. DDS should not be used to oppress the freedoms of citizens or perpetuate government corruption. Critical African philosophers view authoritarianism as governing to accumulate wealth and power rather than serving the needs of their citizens and Africa as a whole (Nyerere, 1962). Even after liberation from colonial rule, some African philosophers accuse their governments of being primarily concerned with replacing the colonial ruling class instead of dismantling it (Coetzee, 2004a; Kohnert, 2022). To maintain their position, government officials focus on maintaining dependent relationships with the West and enforcing cultural nationalism to suppress dissent (Gwagwa, 2019).
@@
African governments have already harnessed their control of national technology through internet shutdowns (Okolo et al., 2023). Therefore, to many authoritarian actors, powerful data technology is just another tool for suppression. Of particular concern to many African practitioners is China as a neocolonial collaborator with African authoritarians. Chinese companies have been found to provide the data technology Ethiopia, Uganda, and Zimbabwe have used to surveil their citizens (Okolo et al., 2023).
@@
While authoritarian uses of data technology are resolutely unethical, more widely accepted uses of government DDS are scrutinized as well. The ubiquitous deployment of a digital ID system forces citizens to choose between access to important services or preserving their privacy from a system they have no control over (Gwagwa, 2019). As governments consider adopting data technology, they need to be accountable to their citizens (Ade-Ibijola and Okonkwo, 2023; Osaghae, 2004). To combat the misuse of government power, DDS should improve government efficiency, transparency, and enforcement of citizens’ freedom (Gwagwa, 2019; Mabe, 2007; Eke et al., 2023a).
4.2. Center All Communities
Community involvement ensures DDS consider the needs and potential impacts of communities beyond the end-users.
Community in Everything. Akan philosophies regard the community as an invaluable resource that guides how every individual lives (Wiredu, 2004a; Metz, 2021; Coetzee, 2004b; Mhlambi and Tiribelli, 2023; Gwagwa et al., 2022). Therefore community input is crucial for constructing a full picture of technical requirements, especially in high-stakes domains (Sinha and Zulfa, 2023; Mhlambi, 2020; Eke et al., 2023a). The concept of community can be misappropriated to deem any collection of stakeholders as sufficient community representatives. African communitarian ethics define a community as individuals with a shared identity who are emotionally invested in each other (Nwankwo and Sonna, 2019; Sinha and Zulfa, 2023; Ruttkamp-Bloem, 2023; Gyekye, 2004). With this more narrow definition of community, involving affected communities in all stages of the lifecycle requires building trust and respecting boundaries by gaining an understanding of cultural norms (Abebe et al., 2021; Ade-Ibijola and Okonkwo, 2023). Additionally, community members should be sufficiently trained or educated on the nature of the technology so they can provide well-informed input (Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023; Adelani et al., 2022; Plantinga et al., 2024). Rather than checking off a list, community-centered data science work should be conducted as a co-creation process in which all stakeholders depend on each other (Langat et al., 2020; Kohnert, 2022; Abebe et al., 2021; Adelani et al., 2022; Nwankwo and Sonna, 2019; Lauer, 2017; Kiemde and Kora, 2022).
Solidarity. Solidarity is understood as looking out for other diverse communities based on mutual respect and the goal of social cohesion (Gwagwa, 2019; Mhlambi, 2020). In Ubuntu understanding, solidarity is a deep care for others, including people of the past, present, future, and the environment (Mhlambi and Tiribelli, 2023; Gwagwa, 2019; Okolo et al., 2023; Dignum, 2023; Gwagwa et al., 2022). With this perspective, DDS should be developed not just with the end user in mind but all the other communities who could be impacted by the technology (Gwagwa et al., 2022; Olojede, 2023; Gyekye, 2004). Solidarity violations between African countries is of particular concern. The success of one African community should not be predicated on the suffering of another (Ndjungu, 2020; Biko, 2004). Upholding solidarity means that all actions made in the data science lifecycle should explicitly protect or improve the lives of vulnerable or marginalized communities.
Exploiting the vulnerability of another is not only unethical but unsustainable due to our interconnected nature. The suffering of one community will eventually lead to the destruction of all communities (Nwankwo and Sonna, 2019). Banding together, “watching one another’s back”, and developing DDS as one big family is key to mitigating harm (Olojede, 2023; Nyerere, 1962).
Inclusion of the Marginalized. While Africa needs to be included in global data science efforts, Africa itself is full of diverse communities that should also be represented in African data science efforts (Adelani et al., 2022; Gwagwa, 2019; Goffi, 2023). African communities’ underrepresentation in datasets across all data science tasks is due to, Gwagwa as described, being uncounted, unaccounted, and discounted (Gwagwa, 2019). Leaving communities out of data also excludes them from the benefits DDS provide (Gwagwa et al., 2022). Given the need to build explicitly African DDS, the lack of African datasets is a threat to efficacy (Okolo et al., 2023; Olojede, 2023; Ade-Ibijola and Okonkwo, 2023). Including marginalized communities requires mutual respect for diverse perspectives and creating procedures such as impact assessments to provide opportunities for inclusive input (African Union, 2024; Abebe et al., 2021; Goffi, 2023; Mhlambi, 2020). In addition, it’s important to challenge the social, political, and economic dynamics that push communities to the margins in the first place (Kiemde and Kora, 2022; Olojede, 2023; Segun, 2021; Day and Rennie, 2023; Uzomah et al., 2023; Abebe et al., 2021).
Center Remote & Rural Communities. Development, especially technical development, is usually focused in urban centers and excludes remote and rural communities (Ade-Ibijola and Okonkwo, 2023; Sinha and Zulfa, 2023; Osaghae, 2004). Given the lack of infrastructure in remote and rural communities, data technology should be used to develop and optimize infrastructures and public services for these regions (Carman and Rosman, 2023; African Union, 2024). However, it’s important to keep in mind that RDS done on behalf of rural and remote communities that do not consider their culture, livelihoods, and direct input can lead to harm (African Union, 2024; Ndjungu, 2020; Carman and Rosman, 2023).
Center Women. Due to the prevalence of patriarchy in many African societies, there is a need to encourage the agency of women in DDS efforts. A few documents suggest that women-led technology businesses and the education of women and girls should be incentivized (African Union, 2024). However, open questions remain about how to maintain African women’s participation in a field known to be male-dominated and antagonistic to women (African Union, 2024; Gwagwa, 2019).
Afro-feminists have a response to the techno-chauvinism that dominates data science (Sinha and Zulfa, 2023). Rather than centering women in general, there must be a recognition of the intersectional status of African women (Sinha and Zulfa, 2023). As articulated by Rosebell Kagumire, African women experience domination through systems of patriarchy, race, sexuality, and global imperialism (Dieng, 2023; Coetzee, 2004b). Therefore, DDS should be developed with the complex needs of African women in mind, because their compounded experiences of marginalization provide insight into the needs of various oppressed populations (Sinha and Zulfa, 2023). There are numerous examples of African women harnessing the internet to fill in the gaps of an oppressive society and data technology holds similar potential (Dieng, 2023). For example, Chil AI Lab Group is a women-led data science collective that is successfully using data technology to address the often neglected health needs of women in Africa (Eke et al., 2023a).
4.3. Uphold Universal Good
Ethical development and deployment of DDS requires a commitment to upholding fundamental human dignity and ensuring these technologies benefit all.
Universal Dignity. Every human and community deserves humane treatment, and DDS should never violate their dignity (Olojede, 2023; Mhlambi, 2020). The African Charter on Human and Peoples’ Rights and the Universal Declaration on Human Rights set precedent for the just treatment of humans (African Union, 2024). Regardless of these laws, African philosophies necessitate respect for human dignity because humans should be inherently valued for their existence and connection to others (Segun, 2021; Metz, 2021; Dignum, 2023). Every human must be treated with respect, care, and concern for their well-being (Wiredu, 2004a; Dieng, 2023; Coetzee, 2004b; Gyekye, 2004; Wiredu, 2004b). In applying the principle of universal dignity to RDS practices, every person involved in the data lifecycle should be respected. Individuals should not be used as a means to execute data work (Metz, 2021; Ramose, 2004b). Rather, all efforts should be taken to ensure their well-being and dignity are preserved when asked to contribute to DDS (Gwagwa, 2019; Abebe et al., 2021). This same respect also extends to communities. Collective agreements need to be honored, and collective work or resources should not be used in a manner that threatens the well-being of the community (Moahi, 2007). While this principle is self-evident, there are many cases in which the rights of Africans were violated for large-scale DDS (African Union, 2024; Segun, 2021; Kohnert, 2022; Moahi, 2007).
Common Good. DDS should contribute to maintaining the safety, health, and goodness of all (Olojede, 2023). In various African philosophies, a person is defined by their commitment to acting for the benefit of those around them (African Union, 2024; Nwankwo and Sonna, 2019; Mhlambi and Tiribelli, 2023; Coetzee, 2004b; Ruttkamp-Bloem, 2023; Gyekye, 2004; Abdul, 2023; Wiredu, 2004b). In terms of RDS, they have to be made with the explicit goal of improving society and dismantling systemic harms (Sinha and Zulfa, 2023; Okolo et al., 2023; Eke et al., 2023a). In Africa, improving the efficacy of agriculture practices, healthcare access, responsiveness of public services, and the security of financial services are over-arching priorities (Carman and Rosman, 2023; Kohnert, 2022). Achieving common good involves incorporating collective values early in the process (Langat et al., 2020; Dignum, 2023), guiding development with regulatory toolkits (Olojede, 2023), not focusing on individualistic profit maximization (Gwagwa, 2019; Segun, 2021; Mabe, 2007; Nyerere, 1962; Mhlambi, 2020; Dieng, 2023), and encouraging the open sharing of data (Gwagwa, 2019; Abebe et al., 2021; Day and Rennie, 2023). Building DDS toward the common good should be the ultimate goal for RDS (Carman and Rosman, 2023; Metz, 2021; Olojede, 2023; Mabe, 2007).
Harmony. DDS should further the mutual well-being of all stakeholders. In addition, data standards and frameworks will be most effective when they harmonize with each other (Gwagwa, 2019; Kiemde and Kora, 2022; Wareham, 2021; Mabe, 2007). In many African philosophies, harmony is not a state but a dynamic and reciprocal process of calibrating one’s actions in response to changes in the environment. In Ubuntu ethics, dogmatism is rejected because it impedes individuals from acting in harmony with the changing world (Ramose, 2004a). In Akan philosophy, morality is defined as acting in line with collective human interests (Wiredu, 2004b). Upholding harmony in data science can be understood on two dimensions: impact and practice. Data should be harnessed to bring people closer to their environment so they can act in the best interests of not only themselves but also those around them. In terms of practice, data ethics frameworks are most effective when all the elements of data science work are accounted for (Kiemde and Kora, 2022; Gwagwa, 2019). Also, acknowledging the unique ethical needs at each stage of the data science lifecycle can inform an adaptable practice of RDS. As Gwagwa et al. assert, the harmonious practice of RDS in Africa requires country-level data ethics frameworks to be in alignment with frameworks developed at the continental level (Gwagwa et al., 2022).
4.4. Communalism in Practice
The development and deployment of DDS should mitigate harms, involve communities in decision-making, and ensure reciprocal benefits for African stakeholders.
Consensus-Building. If data scientists want to develop responsible practices and encourage effective collaboration, consensus-building is a well-practiced strategy from African communities (Wiredu, 2004b). Rather than the majority rule common in Western societies, African elders discuss issues until they all agree on a final decision (Wiredu, 2004a; Carman and Rosman, 2023). Achieving consensus requires the final decision to be 1) the dominant view of the group, 2) in line with the common good, and 3) aligned with the morals of the individual parties (Coetzee, 2004b). Consensus should be broached in an environment of trust, practical reason, humility, openness, and respect for the viewpoints of all involved parties (Coetzee, 2004b; Nwankwo and Sonna, 2019; Gwagwa, 2019; Mhlambi, 2020; Okolo et al., 2023; Gwagwa et al., 2022).
Community engagement provides spaces for consensus in the data science lifecycle to include more perspectives (Day and Rennie, 2023; Mabe, 2007). Consensus processes should also include procedures for documentation to keep track of disagreements, dissenting opinions, and the progression of project values (Kling et al., 2023). It is not easy to achieve these conditions, so conflict management, negotiation, and reasonable bargaining are helpful mechanisms to fully consider and resolve contradicting positions (Gwagwa, 2019; Osaghae, 2004; Dieng, 2023). Consensus should be a dynamic feedback loop to ensure every contributor is on the same page about the team’s approach to RDS (Segun, 2021; Nwankwo and Sonna, 2019; Uzomah et al., 2023; Kohnert, 2022; Abebe et al., 2021; Gwagwa, 2019; Dignum, 2023). The actual process of consensus-building is also a helpful mechanism to build trust between data collaborators and develop informed consent from future users (Nwankwo and Sonna, 2019; Kohnert, 2022; Abebe et al., 2021).
Reciprocity. In many African philosophies, reciprocity is the foundation of a healthy society. In Akan society, practicing reciprocity ensures that community needs are met, while building deep social bonds (Wiredu, 2004b). African perfectionist proponents go as far as to assert that assisting others in achieving their goals makes someone more of a person (Wareham, 2021; Wiredu, 2004b). Without reciprocity, society becomes imbalanced and co-dependent (Coetzee, 2004b; Mhlambi and Tiribelli, 2023; Nyerere, 1962). There are numerous examples of African data subjects not reaping any benefits from the data collaborations they participate in (Gwagwa, 2019; Abebe et al., 2021). This often leads to technically mediated harms while the controllers of data amass profits (Gwagwa, 2019). Therefore, sustainable DDS should practice reciprocity on several dimensions (Wiredu, 2004b). If someone contributes to a DDS they should meaningfully benefit from the system or project (Mhlambi, 2020; Gyekye, 2004; Sinha and Zulfa, 2023). Inspired by philosophies such as Ubuntu or Ujamaa, DDS should operate in a manner that benefits the society in which they are created and deployed (Eke et al., 2023a; Adelani et al., 2022; Dignum, 2023).
Resolving Data-Driven Harms. When disagreements, conflict, or harm occur at any stage of the data science lifecycle, we need mechanisms of accountability and reconciliation to correct wrongs and empower those impacted. In African societies, harm is not just actively making someone’s life worse but also neglecting obligations to the community (Wiredu, 2004a; Gyekye, 2004). A person who causes harm is viewed as a moral failure who must be corrected by their community through sanctions and even mental rehabilitation to correct deeper issues connected to their poor actions (Wiredu, 2004b; Coetzee, 2004b). Even the most powerful members of society, such as chiefs, are subject to correction and even dismissal by their community (Wiredu, 2004a).
The adoption of AI and other DDS have already caused harm to African populations by way of data bias, socio-economic risk, and privacy violations (Ade-Ibijola and Okonkwo, 2023). There are African data ethicists who stress the need to develop procedures for communities and individuals harmed by DDS to seek restitution (Gwagwa, 2019; Mhlambi and Tiribelli, 2023). These solutions are dependent on African governments and external multinational organizations committing to transparency, equality, and restorative practices (Gwagwa, 2019; African Union, 2024; Okolo et al., 2023; Dignum, 2023; Kiemde and Kora, 2022). African governments can mitigate data harm by being transparent about their potential data collaborations, outlining their plans for data protection before, during, and after the deployment of DDS, and enforcing mechanisms of accountability and dissent from their citizens (Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023; Sinha and Zulfa, 2023).
Similar to the dismissal of chiefs, powerful stakeholders acting outside of their agreed duties, must experience restorative consequences, not just a slap on the wrist (Mandaza, 2004; Ndjungu, 2020; Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023; Mhlambi and Tiribelli, 2023; Langat et al., 2020; Gwagwa, 2019; Mhlambi, 2020; Biko, 2004; Coetzee, 2004a).
Fair Collaboration. Given the current gap between Africa’s AI readiness and growing interest in AI adoption, many concede external partnership as a necessity(Eke et al., 2023b; African Union, 2024). However, exploitative external relationships set a precedent that curtails African self-determination in data science work (Ndjungu, 2020; Sinha and Zulfa, 2023). When building relationships, there are established obligations that each collaborator owes to the other (Coetzee, 2004b; Metz, 2021). Data collaborations need to be predicated on trust, fair attribution of work, and a commitment to prioritizing the agency of African collaborators (Adelani et al., 2022; Nwankwo and Sonna, 2019; Abebe et al., 2021; Gwagwa, 2019; Wareham, 2021; Wiredu, 2004b).
4.5. Data Self-Determination

### v3

@@ 4. African Data Science Ethical Framework
In this section, we summarize the key components of our framework for African data ethics. For an overview of the six major principles, 19 associated minor principles and recommended practices or policies derived from our analysis, view Table 4.
| Major Principle | Minor Principle | Practices & Policies |
| Challenge Power Asymmetries | Challenge Colonial Power | Adopt data science practices that are localized, decentralized, bottom-up and in less deference to Big Tech paradigms |
| Challenge Internal Power Asymmetry | Establish agencies to triage citizen concerns and doubts about new technology | |
| Assert Data Self-Determination | ”For Africans, By Africans” | Africans speak for themselves in global data ethics discourses, negotiations for technical partnerships and design for local needs |
| Treasure Indigenous Knowledge | Conduct natural language processing projects to document and represent indigenous languages | |
| Protect Data Agency | Integrate communal conceptions of privacy for data protection such as data trusts, data collaboration tools, and community data management | |
| Invest in Local Data Institutions & Infrastructures | Invest in Physical & Organizational Infrastructures | Construct secure and locally-run peer to peer networks |
| Cultivate Governance Infrastructure | Develop protocol for conducting algorithmic impact assessments | |
| Support Formal & Informal Data Science Collectives | Provide technical training and education for data scientists outside of the university and working in the informal sector | |
| Utilize Communalist Practices | Ground Design in Community-Engaged Consensus | Consult intended users on a routine basis to build trust and incorporate their cultural perspectives in design |
| Instantiate Reciprocal Relationships | Collaborate with African organizations already conducting research or building technologies of interest | |
| Mend Data-Driven Harms | Implement reconciliation processes that place disenfranchised communities in positions of power and hold the perpetrating parties accountable | |
| Center Communities on the Margins | Design in Solidarity | Embrace diverse data ethics standards and avoid imposing standards on other communities |
| Center Remote & Rural Communities | Develop AI solutions that address the needs of remote communities such as agriculture, climate change, and healthcare accessibility | |
| Center Women | Incentivize women-led data science entrepreneurship | |
| Protect & Empower Youth | Organize continental meetings for young Africans to discuss the state of data science in their communities | |
| Uphold Common Good | Adopt Technology with a Measured Mindset | Appraise imported state-of-art technologies before use |
| Preserve the Dignity of Data Contributors | Treat every collaborator as morally responsible agents deserving of materially and psychologically safe working conditions | |
| Strive for Common Good with Systemic Change | Build open data ecosystems that safeguard against harmful use and compensate data subjects from economic gains | |
| Maintain Harmony with Stakeholders | Update country-level data ethics to align with continental frameworks and vice versa |
A table of the major, minor and practices/policies surfaced from our African Data Ethics corpus. The ordering of major and minor principles follows the order in section 4.
4.1. Challenge Power Asymmetries
@@
Challenge Colonial Power. RDS practices from the West do not seamlessly transfer to the African context because these practices are developed within colonial contexts disconnected from the realities of African practitioners and users (Eke and Ogoh, 2022; Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023; Adelani et al., 2022; Gwagwa, 2019; Eke et al., 2023b; Okolo et al., 2023; Eke et al., 2023a; Goffi, 2023; Carman and Rosman, 2023). African practitioners identify three dimensions in which the sociotechnical disconnect harms responsible data science in Africa: epistemic injustice, dehumanizing extraction, and dependent partnerships. Firstly, African scholars identify how philosophical epistemic injustice permeates global data ethics paradigms (Eke and Ogoh, 2022; Metz, 2021; Olojede, 2023). As many African philosophers agree, Enlightenment ideals (a premier part of the Western philosophical canon) were predicated on colonialism and anti-black racism (Gwagwa, 2019). European colonizers weaponized the Enlightenment emphasis on human rationality to deem Africans sub-human by assessing them as primitive and incapable of reason (Lauer, 2017). Colonization was not only justified but encouraged by the philosophies of rationality. Colonizers determined they had a moral obligation to develop Africans capacity for reason through European education. The legacy of colonialism is why African data scientists encourage casting aside Western perspectives to develop distinctly African RDS perspectives (Mhlambi, 2020). Additionally, an over reliance on “rational” performance metrics encourages the same epistemic injustice that justifies and encourages the marginalization of Africans in technology such as facial recognition (Mhlambi, 2020; Buolamwini and Gebru, 2018; Cisse, 2018; Gwagwa et al., 2022).
@@
Secondly, many documents recognize most African contributions to data science as disproportionately benefiting corporations such as OpenAI, Google, Meta, and Microsoft (Ndjungu, 2020; Chan et al., 2021; Abebe et al., 2021; Nwankwo and Sonna, 2019; Kiemde and Kora, 2022). The computing demand of large-data systems such as AI proliferates neocolonialism to new heights in Africa (Eke et al., 2023a). The work of Africans within the data science ecosystem should benefit Africans first (Kohnert, 2022). The fact that it currently does not is connected to the legacy of colonialism and chattel slavery in which Africans were forced to extract their raw materials so colonialists could fuel industrialization and capitalism in their home countries (Mhlambi, 2020; Ndjungu, 2020; Day and Rennie, 2023; Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023; Birhane, 2020).
@@
Finally, the last vestige of colonial power to be challenged in African RDS are dependent partnerships. Africa currently lacks the technical infrastructure for large-scale data-driven technology, which pushes data scientists towards unfair agreements with powerful organizations to access vital resources (Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023; Hountondji, 2004; Osaghae, 2004). Even worse, companies such as Amazon, Google, Meta, and Uber use savior language such as “liberating the bottom million” to describe their digital services in Africa (Abebe et al., 2021).
@@
Challenge Internal Power Asymmetry. Critical African philosophers conceptualize authoritarianism as governing to accumulate wealth and power rather than serving the needs of citizens and Africa as a whole (Nyerere, 1962). Even after liberation from colonial rule, many African philosophers accuse their governments of replacing the colonial ruling class instead of dismantling it (Coetzee, 2004a; Kohnert, 2022). To maintain their position, government officials focus on maintaining dependent relationships with the West and enforcing cultural nationalism to suppress dissent (Gwagwa, 2019).
@@
African governments have already harnessed their control of national technology through internet shutdowns (Okolo et al., 2023). Therefore, to many authoritarian actors, powerful data technology is just another tool for suppression. Of particular concern to many African practitioners is China as a neocolonial collaborator with African authoritarian leaders and governments. Chinese companies have been found to provide the data technology Ethiopia, Uganda, and Zimbabwe have used to surveil their citizens (Okolo et al., 2023).
@@
While authoritarian uses of data technology are resolutely unethical, more widely accepted uses of government data-driven technologies are scrutinized as well. The ubiquitous deployment of digital identification systems force citizens to choose between accessing important services or preserving their privacy from a system they have no control over (Gwagwa, 2019). As governments consider adopting data technology, they need to be held accountable to their citizens (Ade-Ibijola and Okonkwo, 2023; Osaghae, 2004). To combat the misuse of government power, authors in our corpus suggest data science initiatives should focus on improving government efficiency, transparency, and enforcement of citizens’ freedom (Gwagwa, 2019; Mabe, 2007; Eke et al., 2023a).
4.2. Assert Data Self-Determination